# PixArt-Σ: Weak-to-Strong Training of Diffusion Transformer for 4K Text-to-Image Generation

### v1

@@ 1 Introduction
To explore this issue, our research focuses on enhancing PixArt-$\alpha$, an efficient T2I training method. PixArt-$\alpha$ represents an early venture within the DiT framework, a model structure with significant potential, as evidenced by works such as Sora [32] and Stable Diffusion 3 [41]. To maximize this potential, we build upon the pre-trained foundation of PixArt-$\alpha$, integrating advanced elements to facilitate its continuous improvement, resulting in a more powerful model, PixArt-Σ. We refer to this process of evolving from a relatively weaker baseline to a stronger model through efficient training as “weak-to-strong training”. Specifically, to achieve “weak-to-strong training”, we introduce the following enhancements:
@@ 2 Related Work
Diffusion Transformers. The Transformer architecture has achieved remarkable success across various domains, such as language modeling [36, 37], computer vision[42, 22, 50, 52], and other areas [13, 3]. In the realm of diffusion models, DiT [33] and UViT [2] pioneer the use of Transformer architecture. Subsequent works, including DiffiT [14], SiT [27], and FiT [25], have improved upon DiT’s architecture, while  [11, 51] enhance training efficiency through masked modeling techniques. For Text-to-Image (T2I) synthesis, PixArt-$\alpha$ [4] explore efficient T2I training schemes, achieving the first Transformer-based T2I model capable of generating 1024px high-quality images. The recent advent of the powerful video generation model Sora [32] has further underscored the potential of Diffusion Transformers. In this work, for the first time, we explore using the Transformer architecture to generate 4K ultra-high-resolution images directly, tackling the computational complexity challenges posed by involving long-sequence tokens.
@@ 3.1 Data Analysis
Interestingly, we have observed that as the resolution of the images increases, there is an improvement in model’s fidelity (Fréchet Inception Distance (FID) [16]) and semantic alignment (CLIP Score), which underscores the importance of the capabilities of generating high-resolution images.
@@ Appendix 0.A Appendix
0.A.1 Training Details
In Tab. 4, we provide detailed information on each training stage of PixArt-Σ, including image resolution, the total volume of training samples, the number of training steps, batch size, learning rate, and the computing time measured in GPU days. Utilizing the Internal-Σ dataset and integrating a more advanced VAE, our proposed method quickly adapts to the new VAE using only 5 V100 GPU days. Subsequently, we only use 50 V100 GPU days to attain a superior text-image alignment.
Notably, applying KV token compression stands out as a significant efficiency booster, markedly reducing the training duration. For instance, fine-tuning from 512px to 1024px and implementing KV compression decreases the required training time from 50 V100 GPU days to just 20 V100 GPU days. Similarly, for resolutions of 2K and 4K, the training time is reduced from 10 to 7 A800 GPU days and 16 to 12 A800 GPU days, respectively. This showcases the efficacy of KV token compression in enhancing training efficiency.
Table 4: We report detailed information about every PixArt-Σ training stage in our paper. Note that Internal-Σ dataset here includes 33M internal data. The count of GPU days excludes the time for VAE feature extraction and T5 text feature extraction, as we offline prepare both features in advance so that they are not part of the training process and contribute no extra time to it.
@@ 0.A.1 Training Details
0.A.2 PixArt-Σ vs. T2I products
@@
0.A.3 More images generated by PixArt-Σ
Fig. 12,  14, and  13 showcase additional visual outputs produced by PixArt-Σ. The quality of these samples is remarkable, characterized by their high fidelity and accuracy in closely matching the provided textual prompts.
@@ 0.A.3 More images generated by PixArt-Σ
Figure 12: Illustrations of High-quality images generated by PixArt-Σ. PixArt-Σ is able to generating high-quality images with fine-grained details, and diverse images with different aspect ratios.

### v2

@@ 1 Introduction
To explore this issue, our research focuses on enhancing PixArt-$\alpha$, an efficient T2I training method. PixArt-$\alpha$ represents an early venture within the DiT framework, a model structure with significant potential, as evidenced by works such as GenTron [7], Sora [34] and Stable Diffusion 3 [43]. To maximize this potential, we build upon the pre-trained foundation of PixArt-$\alpha$, integrating advanced elements to facilitate its continuous improvement, resulting in a more powerful model, PixArt-Σ. We refer to this process of evolving from a relatively weaker baseline to a stronger model through efficient training as “weak-to-strong training”. Specifically, to achieve “weak-to-strong training”, we introduce the following enhancements:
@@ 2 Related Work
Diffusion Transformers. The Transformer architecture has achieved remarkable success across various domains, such as language modeling [38, 39], computer vision[44, 24, 53, 55], and other areas [15, 3]. In the realm of diffusion models, DiT [35] and UViT [2] pioneer the use of Transformer architecture. Subsequent works, including DiffiT [16], SiT [29], and FiT [27], have improved upon DiT’s architecture, while  [13, 54] enhance training efficiency through masked modeling techniques. For Text-to-Image (T2I) synthesis, PixArt-$\alpha$ [5] explore efficient T2I training schemes, achieving the first Transformer-based T2I model capable of generating 1024px high-quality images. GenTron [7] explores the flexibility and scalability of diffusion Transformers in both image and video generation. The recent advent of the powerful video generation model Sora [34] has further underscored the potential of Diffusion Transformers. In this work, for the first time, we explore using the Transformer architecture to generate 4K ultra-high-resolution images directly, tackling the computational complexity challenges posed by involving long-sequence tokens.
@@ 3.1 Data Analysis
Interestingly, we have observed that as the resolution of the images increases, there is an improvement in the model’s fidelity (Fréchet Inception Distance (FID) [18]) and semantic alignment (CLIP Score), which underscores the importance of the capabilities of generating high-resolution images.
@@ 6 Conclusion
Acknowledgement. We would like to thank Zeqiang Lai, Fei Chen, and Shuchen Xue for discussing and helping implement the DMD part.
@@ Appendix 0.A Appendix
0.A.1 Extension: Inference Acceleration (PixArt + DMD).
Experiments. To expedite the inference, we integrate PixArt with DMD [52], a one-step inference technique achieved through distribution matching distillation. We train a one-step Generator $G_{\theta}$, and the generated image is denoted as $x_{0}=G_{\theta}(\overline{T},\rm{text})$. Initially, we set $\overline{T}$ to the same value as the denoising timesteps during training, that is $\overline{T}=999$ following [52]. However, we observed undesired results and investigated the most suitable value for $\overline{T}$ as shown in Fig. 10. Surprisingly, the optimal $\overline{T}$ was found to be 400 rather than 999. This deviation arises from the fact that a smaller $T$ enhances the model’s confidence in predicting the noise of $G_{\theta}$. However, this principle is effective only within a certain range. If $T$ becomes too small, the scenario significantly deviates from the training setting of the base model. Therefore, there exists a trade-off in selecting the $T$ value. Finally, we compare our method quantitatively as well as qualitatively with PixArt + LCM [4] in Tab. 4 and Fig. 11. We also provide a demo video (attached as dmd_demo.mp4) to compare the speed of PixArt + DMD with the teacher model.
Figure 10: Base model + DMD performance over $\overline{T}$.
Table 4: Comparison of PixArt + DMD performance compared to PixArt + LCM. These experiments are conducted on 512x512 resolution with a batch size of 1.
[…]
PixArt + LCM (1 step) 108.66 0.2247 0.11s
PixArt + LCM (2 step) 17.95 0.2736 0.16s
PixArt + LCM (4 step) 13.06 0.2797 0.26s
PixArt + DMD (1 step) 13.35 0.2788 0.11s
Teacher model (20 steps) 9.273 0.2863 1.44s
Figure 11: Visual comparison of Base model + DMD performance vs. Base model + LCM model.
0.A.2 Training Details of PixArt-Σ
In Tab. 5, we provide detailed information on each training stage of PixArt-Σ, including image resolution, the total volume of training samples, the number of training steps, batch size, learning rate, and the computing time measured in GPU days. Utilizing the Internal-Σ dataset and integrating a more advanced VAE, our proposed method quickly adapts to the new VAE, requiring a mere 5 V100 GPU days. Subsequently, we achieve remarkable text-image alignment with just 50 V100 GPU days, showcasing a substantial improvement over PixArt-$\alpha$ at minimal additional training expense, thanks to the proposed "weak-to-strong" strategy, which proves highly efficient.
Notably, applying KV token compression stands out as a significant efficiency booster, markedly reducing the training duration. For example, when fine-tuning from 512px to 1024px and incorporating KV compression, the training time reduces dramatically from 50 V100 GPU days to a mere 20 V100 GPU days. Likewise, for resolutions of 2K and 4K, the training time diminishes from 20 to 14 A800 GPU days and from 25 to 20 A800 GPU days, respectively. This underscores the remarkable efficacy of KV token compression in augmenting training efficiency.
Table 5: We report detailed information about each training stage of PixArt-Σ. Note that Internal-Σ dataset here includes 33M internal data. The count of GPU days excludes the time for VAE feature extraction and T5 text feature extraction, as we offline prepare both features in advance so that they are not part of the training process and contribute no extra time to it.
@@ 0.A.2 Training Details of PixArt-Σ
0.A.3 Detailed Settings of the Ablation Studies
In this subsection, we describe the additional experimental setup for the ablation studies. For each study, we evaluated the Fréchet Inception Distance (FID) by comparing images generated by our PixArt-Σ using a set of 30,000 curated prompts against those from the High-Quality Evaluation Dataset. Additionally, we assessed the Clip-Scores by comparing the model-generated images to the original 30,000 prompts. Our findings suggest that the number of images used for testing can influence the FID scores; for instance, a dataset comprising 10,000 images typically yields higher FID scores. Besides, for the conducted experiments on compression positions and operators, we test the respective models with images generated at a resolution of 512px.
0.A.4 Samples of High-Quality Evaluation Dataset
Figure 12: Samples in our proposed High-Quality Evaluation Dataset. The evaluation dataset presented in this paper contains samples of superior visual quality compared to those in COCO-30K.
We observe the evaluations performed on the MSCOCO dataset could not adequately fully capture a model’s proficiency in aesthetics and text-image alignment [40, 21]. Thus, we propose an evaluation dataset that consisting of 30,000 high-quality, aesthetically pleasing text-image pairs for a more thorough assessment of a model’s ability to generate visually appealing images. The visualizations in Fig. 12 exemplify the dataset’s superior aesthetic quality and the precise alignment between textual descriptions and visual content.
0.A.5 FID Comparisons with Open-source Models
Table 6: Comparisons on FID and Clip-Score with Open-sourced T2I Models. PixArt-Σ demonstrates enhanced performance in terms of FID and Clip-Score on the curated 30K High-Quality Evaluation Dataset.
[…]
Stable 1.5 0.9 17.03 0.2748
Stable Turbo 3.1 10.91 0.2804
Stable XL 2.6 7.38 0.2913
Stable Cascade 5.1 9.96 0.2839
Playground-V2.0 2.6 8.68 0.2885
Playground-V2.5 2.6 7.64 0.2871
PixArt-$\alpha$ 0.6 8.65 0.2787
PixArt-Σ 0.6 8.23 0.2797
We conducted comparative analyses of open-source models using FID and CLIP-Scores on our curated dataset. The results, presented in Tab. 6, reveal that the weak-to-strong fine-tuning process significantly improves the model’s ability to generate high-quality images and to align more closely with the given instructions, as PixArt-Σ shows a lower FID (8.65 v.s. 8.23) and higher Clip-Scores (0.2787 v.s 0.2797) compared to PixArt-$\alpha$. Besides, compared to other models, the PixArt-Σ still shows comparable or even better performance with relatively small network parameters (0.6B).
0.A.6 Designed Prompts for AI preference Study
To study AI preferences among various text-to-image (T2I) generators, we employ the advanced multi-modality model, GPT-4V, as an automated evaluator. The instructions given to GPT-4V for comparing the generated images based on quality and adherence to prompts are illustrated in Fig. 13. GPT-4V demonstrates the ability to provide logical assessments and reasons that coincide with human preferences.
Figure 13: Illustration of the designed prompts for GPT-4Vision.: We developed various prompts for GPT-4V to assess the image quality and instruction-following ability of different T2I generators, respectively.
0.A.7 Pseudo-code for KV Token Compression
We present the PyTorch-style pseudo-code for the KV Token Compression algorithm described in Sec.3.2 in the main paper. The implementation is straightforward.
Algorithm 1 KV Token Compression.
import torch
import torch.nn as $n$
import xformer
class AttentionKVCompress($n$.Module):
def __init__(self, dim, sampling=’conv’, sr_ratio=1,, **kwargs):
super().__init__()
# Projection layers and non-relavent definitions are omitted.
self.sampling=sampling # [‘conv’, ‘ave’, ‘uniform’]
self.sr_ratio = sr_ratio
if sr_ratio > 1 and sampling == "conv":
# Avg Conv Init.
self.sr = $n$.Conv2d(dim, dim, groups=dim, kernel_size=sr_ratio, stride=sr_ratio)
self.sr.weight.data.fill_(1/sr_ratio**2)
self.sr.bias.data.zero_()
self.norm = $n$.LayerNorm(dim)
def downsample_2d(self, tensor, H, W, scale_factor, sampling=None):
B, N, C = tensor.shape
tensor = tensor.reshape(B, H, W, C).permute(0, 3, 1, 2)
new_H, new_W = int(H / scale_factor), int(W / scale_factor)
new_N = new_H * new_W
if sampling == "ave":
tensor = F.interpolate(
tensor, scale_factor=1 / scale_factor, mode=’nearest’
).permute(0, 2, 3, 1)
elif sampling == "uniform":
tensor = tensor[:, :, ::scale_factor, ::scale_factor].permute(0, 2, 3, 1)
elif sampling == "conv":
tensor = self.sr(tensor).reshape(B, C, -1).permute(0, 2, 1)
tensor = self.norm(tensor)
return tensor.reshape(B, new_N, C).contiguous(), new_N
def forward(self, x, mask=None, HW=None, block_id=None, ):
B, N, C = x.shape
new_N = N
H, W = HW
qkv = self.qkv(x).reshape(B, N, 3, C)
q, k, v = qkv.unbind(2)
if self.sr_ratio > 1:
k, new_N = self.downsample_2d(k, H, W, self.sr_ratio, sampling=self.sampling)
v, new_N = self.downsample_2d(v, H, W, self.sr_ratio, sampling=self.sampling)
x = xformers.ops.memory_efficient_attention(q, k, v)
x = x.view(B, N, C)
x = self.proj(x)
x = self.proj_drop(x)
return x
0.A.8 PixArt-Σ vs. T2I products
@@ 0.A.7 Pseudo-code for KV Token Compression
0.A.9 More images generated by PixArt-Σ
Fig. 16,  17 and  18 showcase additional visual outputs produced by PixArt-Σ. The quality of these samples is remarkable, characterized by their high fidelity and accuracy in closely matching the provided textual prompts.
0.A.10 Limitation and Social Impact
Limitation. Our model still lacks the ability to generate some specific scenes and objects, especially text generation and hand generation. It is not perfect in the following aspects: it cannot fully align the complex prompts input by the user, face generation may have flaws and sensitive content may be generated. Subsequent research should focus on higher-quality data construction, scale model size, and improving model illusion and security issues through super alignment.
Negative social impact. Text-to-image models may bring a negative social impact by generating images that present stereotypes or discriminate against certain groups. For instance, the images generated by a text-to-image model may depict unbalanced proportions of gender and inaccurate content for some uncommonly used concepts. Mitigating these issues requires careful data collection.
@@ 0.A.10 Limitation and Social Impact
Figure 16: Illustrations of High-quality images generated by PixArt-Σ. PixArt-Σ can generate high-quality images with fine-grained details, and diverse images with different aspect ratios.